OpenStudy (swaqqout_kid1):

Seeley is using a cart to move several large boxes at a time. If each box weighs 22 lb and the cart weighs 11 lb, how much total weight must he move if he loads 5 boxes onto the cart? 121 lb 165 lb 77 lb 110 lb @sangya21

OpenStudy (anonymous):

total weight = 22(x) + 11 for x = 5 w = 22(5) + 11